Saxbury Corporations relevant range of actvity is 3,000 units to 7000 units. When it produces and sells 4 per unit are as follows: Cont pex hit 6.80 s 3.50 1.60 4.00 0.75 0.60 0.50 0.50 Direct naterials Direct labor 0256:22 Variable manufacturing overhead Pixed manufacturing overhead Pixed selling expense Pixed administrative expense Sales commissioas Variable administrative expense Required: a. For financial reporting purposes, what is the total amount of product costs incurred to make 4,000 units? b. For financial reporting purposes, what is the total amount of period costs incurred to sell 4,000 units? c. If 5,000 units are sold, what is the variable cost per unit sold? (Round Per unit answer to 2 decimal places) d. If 5,000 units are sold, what is the total amount of variable costs related to the units sold? e. If 5,000 units are produced, what is the average fixed manufacturing cost per unit produced? (Round Per decimal places.) f. If 5,000 units are produced, what is the total amount of fixed manufacturing cost incurred? g. If 5,000 units are produced, what is the total amount of manufacturing overhead cost incurred? What is this total amount expressed on a per unit basis? (Round Per unit answer to 2 decimal places.) h. If the selling price is $24.30 per unit, what is the contribution margin per unit sold? (Round Perunit answer to 2 decimal places.) L, If 3,000 units are produced, what is the total amount of direct manufacturing cost incurred? J. If 3,000 units are produced, what is the total amount of indirect manufacturing cost incurred? k. What incremental manufacturing cost will the company incur if it increases production from 4.000 to 4,001 units? (Round Per unit answer to 2 decimal places)
media%2F73f%2F73fd3e75-017b-437a-a624-cb
Answer #1

a) Total product Cost = (6.8+3.5+1.6+4)*4000 = $63600

b) Total period cost = (0.75+.60+.50+.50)*4000 = $9400

c) Variable cost per unit sold = 6.8+3.5+1.6+.5+.5 = $12.90 per unit

d) Total variable cost = 12.9*4000 = $51600

Note; Please post whole question in the group of 4-4 parts..